Find all the fixed points and classify each of them: f(x, y) = xye−x
2−y
2
.
1
Expert's answer
2021-08-05T13:58:58-0400
"f(x, y)=xye^{-x^2-y^2}"

"f_x=ye^{-x^2-y^2}-2x^2ye^{-x^2-y^2}"

"f_y=xe^{-x^2-y^2}-2xy^2e^{-x^2-y^2}"

"\begin{matrix}
 f_x=0 \\
 f_y=0
\end{matrix}=>\begin{matrix}
 ye^{-x^2-y^2}-2x^2ye^{-x^2-y^2}=0 \\
 xe^{-x^2-y^2}-2xy^2e^{-x^2-y^2}=0
\end{matrix}"

"\begin{matrix}
 y(1-2x^2)=0 \\
 x(1-2y^2)=0
\end{matrix}"

Critical points:

"(-\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2}), (-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2}), (0, 0),"

"(\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2}), (\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})"

"f_{xx}=-6xye^{-x^2-y^2}+4x^3ye^{-x^2-y^2}"

"f_{xy}=e^{-x^2-y^2}-2y^2e^{-x^2-y^2}-2x^2e^{-x^2-y^2}+4x^2y^2e^{-x^2-y^2}"

"f_{yy}=-6xye^{-x^2-y^2}+4xy^3e^{-x^2-y^2}"

"(-\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})"


"f_{xx}=-2e^{-1}<0"

"f_{xy}=0"

"f_{yy}=-2e^{-1}"

"\begin{vmatrix}
 -2e^{-1} & 0 \\
 0 & -2e^{-1}
\end{vmatrix}=4e^{-2}>0"

Point "(-\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})" is a local maximum.



"(-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})"


"f_{xx}=2e^{-1}>0"

"f_{xy}=0"

"f_{yy}=2e^{-1}"

"\begin{vmatrix}
 2e^{-1} & 0 \\
 0 & 2e^{-1}
\end{vmatrix}=4e^{-2}>0"

Point "(-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})" is a local minimum.



"(\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})"

"f_{xx}=2e^{-1}>0"

"f_{xy}=0"

"f_{yy}=2e^{-1}"

"\begin{vmatrix}
 2e^{-1} & 0 \\
 0 & 2e^{-1}
\end{vmatrix}=4e^{-2}>0"

Point "(\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})" is a local minimum.


"(\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})"


"f_{xx}=-2e^{-1}<0"

"f_{xy}=0"

"f_{yy}=-2e^{-1}"

"\begin{vmatrix}
 -2e^{-1} & 0 \\
 0 & -2e^{-1}
\end{vmatrix}=4e^{-2}>0"

Point "(\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})" is a local maximum.



"(0,0)"


"f_{xx}=0"

"f_{xy}=1"

"f_{yy}=0"

"\begin{vmatrix}
 0 & 1\\
 1 & 0
\end{vmatrix}=-1<0"

Point "(0,0)" is a saddle point.